OpenStudy (sstradic):

The legs of a right triangle are 5 units and 15 units. What is the length of the hypotenuse

OpenStudy (sstradic):

steve help

OpenStudy (steve816):

Use the Pythagorean theorem! \[\large a^2+b^2=c^2\]

OpenStudy (steve816):

Where c is the hypotenuse and a and b are the legs

OpenStudy (steve816):

So then it would be\[\large 5^2+15^2=c^2\] And then, just solve for c and you will get the answer!

OpenStudy (steve816):

Got it?

OpenStudy (sstradic):

so radical 250 or radical 200 :/ I'm stuck

OpenStudy (steve816):

square root of 250 it is.

OpenStudy (sstradic):

yay i told my mom i wasent autistic but she didn't me believe no me big boy answer question by self hah mom face in your lel